# Linear-quadratic optimal control for abstract differential-algebraic equations

### v1

@@ 3 Exponential stability of abstract differential-algebraic equations
Utilizing the generator $A_{R}$ as described in (17), we can extend the characterization of exponential stability from Proposition 2.3 to pseudo-resolvents and degenerate semigroups. In this context, for a given degenerate semigroup $(T(t))_{t\geq 0}$ that is exponentially stable in the sense that it fulfills (19), we apply Laplace transform yielding a pseudo resolvent that is defined for all $\lambda\in\mathbb{C}$ with $\operatorname{Re}(\lambda)>\omega$, as discussed in [Are01], namely
@@
    There exists a positive operator $P\in L(X)$ such that
    $\langle R(0)x,Px\rangle+\langle Px,R(0)x\rangle=-\langle R(0)x,R(0)x\rangle,\quad$
@@
    There exists a positive operator $P\in L(X)$ such that
    $\langle R(0)x,Px\rangle+\langle Px,R(0)x\rangle\leq-\langle R(0)x,R(0)x\rangle,\quad$
@@
    $(-\omega,\infty)\subseteq\rho(E,A)$ and $(-\omega,\infty)\subseteq\rho(E^{*},A^{*})$
@@
In particular, the degenerated semigroup associated with the pseudo-resolvent $\lambda\mapsto(A-\lambda E-A)^{-1}E$ is exponentially stable.
@@ 4 Linear quadratic optimal control of abstract differential-algebraic equations
We consider the pseudo-resolvent $R_{r}$ as in (14), and we assume that it fulfills the growth condition (15) for $\Omega=\{\lambda\in\mathbb{C}~{}|~{}\operatorname{Re}\lambda>\omega\}$ for some $\omega<0$. Further, let $T$ be the associated exponentially stable degenerated semi-group. With the notation as in Proposition 2.5, we use the bounded projection $P$ which fulfills ${\rm ran}\,P=X_{R}$ and $\ker P=X_{K}$ and decompose $A^{-1}Bu=(I-P)A^{-1}Bu+PA^{-1}Bu=x_{0}+x_{1}\in X$ and set $\hat{B}_{0}u:=(I-P)A^{-1}Bu$ and $\hat{B}_{1}u:=PA^{-1}Bu$ for all $u\in U$. Then $\hat{B}_{1}\in L(U,X_{R})$, $\hat{B}_{0}\in L(U,X_{K})$ and the mild solution of (11) is given by
@@
$\Psi:\quad X\rightarrow L^{2}([0,t_{f}],Y),\quad(\Psi x_{0})(t):=CT(t)x_{0},\quad\text{$t\in[0,t_{f}]$}.$
By invoking the boundedness assumption (10), we can, by using [Wei91], infer that the input-output operator
@@
$y=\Psi x_{0}+\mathbb{F}u.$ (29)
In the following we consider the time-varying weights $\mathcal{Q}(\cdot)\in L^{\infty}([0,t_{f}],\mathbb{R}^{n\times n})$, $\mathcal{R}(\cdot)\in L^{\infty}([0,t_{f}],\mathbb{R}^{m\times m})$, which are assumed to be symmetric almost everywhere and $\mathcal{N}(\cdot)\in L^{\infty}([0,t_{f}],\mathbb{R}^{m\times n})$ which are used in the quadratic cost function
$\hat{J}(u,y)=\int_{0}^{t_{f}}\left\langle\begin{pmatrix}y(t)\\ u(t)\end{pmatrix},\begin{pmatrix}\mathcal{Q}(t)&\mathcal{N}^{*}(t)\\ \mathcal{N}(t)&\mathcal{R}(t)\end{pmatrix}\begin{pmatrix}y(t)\\ u(t)\end{pmatrix}\right\rangle_{Y\times U}{\rm d}t=\left\langle\begin{pmatrix}y\\ u\end{pmatrix},\begin{pmatrix}\mathcal{Q}&\mathcal{N}^{*}\\ \mathcal{N}&\mathcal{R}\end{pmatrix}\begin{pmatrix}y\\ u\end{pmatrix}\right\rangle_{L^{2}}.$
@@
$J(u,x_{0}):=\hat{J}(u,\Psi x_{0}+\mathbb{F}u)=\left\langle\begin{pmatrix}x_{0}\\ u\end{pmatrix},\begin{pmatrix}\Psi^{*}\mathcal{Q}\Psi&\Psi^{*}(\mathcal{Q}\mathbb{F}+\mathcal{N}^{*})\\ (\mathbb{F}^{*}\mathcal{Q}+\mathcal{N})\Psi&\mathscr{R}\end{pmatrix}\begin{pmatrix}x_{0}\\ u\end{pmatrix}\right\rangle$
@@
Assume that we have an abstract differential-algebraic system satisfying (15) and that $\mathscr{R}$ satisfies the coercivity assumption (31). Then for all $x_{0}\in X$
@@
$u^{\rm opt}(t)=-\mathscr{R}^{-1}(\mathbb{F}^{*}\mathcal{Q}+\mathcal{N})\Psi(x_{0})(t),\quad\text{for a.e.\ $t\geq 0$}.$ (32)
@@
We use an idea from [WW97], abbreviate $\hat{u}:=u+\mathscr{R}^{-1}(\mathbb{F}^{*}\mathcal{Q}+\mathcal{N})\Psi(x_{0})$ and complete the square as follows
@@
Since $\mathscr{R}$ is uniformly positive the minimum is attained if and only if the second term vanishes. ∎
The formula (32) for the unique minimizer can be rewritten with $y(t)=\Psi(x_{0})(t)+\mathbb{F}u_{opt}(t)$ as
@@
In addition to the assumptions of Proposition 4.1 let the cost function $J$ satisfy $\mathcal{N}=0$ and let there exists $\varepsilon>0$ such that $\langle\mathcal{R}u,u\rangle\geq\varepsilon\|u\|$ holds for all $u\in L^{2}([0,t_{f}],U)$. Then
$u^{\rm opt}(t)=-(I-\mathscr{R}^{-1}\mathbb{F}^{*}\mathcal{Q}\mathbb{F})^{-1}\mathscr{R}^{-1}\mathbb{F}^{*}\mathcal{Q}y(t).$
@@
$\|\mathscr{R}^{-1}\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|\leq\frac{\|\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|}{\|\mathcal{R}+\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|}\leq\frac{\|\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|}{\varepsilon+\|\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|}<1$
therefore the inverse of $I-\mathscr{R}^{-1}\mathbb{F}^{*}\mathcal{Q}\mathbb{F}$ can be obtained from the Neumann series. ∎
If the coercivity condition (31) is violated then the existence of an optimal control input $u_{opt}$ is not guaranteed, which can already be observed for finite-dimensional DAEs [RV19].
Remark 4.3.
@@
In particular, the result of Proposition 4.1 and Corollary 4.2 continues to hold. Well-posed linear systems that satisfy (33) with $E=Id$ are called externally stable in the literature. For ODEs it was shown that for exponentially stabilizable and detectable systems the conditions (33) are equivalent to the exponential stability of the system (also called internal stability), see [Reb93].
@@
The mild solutions that are considered here can be characterized using the mild solutions of the pseudo-resolvent equation (21) with $R=R_{r}$. It was shown in [GR23, Lemma 8.1] that one can also rewrite the ADAE (11) using the pseudo-resolvent $R=R_{l}=EA^{-1}$.
$\frac{{\rm d}}{{\rm d}t}EA^{-1}z(t)=z(t)+f(t),\quad EA^{-1}z(0)=EA^{-1}z_{0}$ (35)
which does not change the inhomogeneity, but to guarantee the existence of solutions the initial value of the DAE is restricted to $x_{0}=A^{-1}z_{0}\in D(A)$.
@@
    $J(u_{\omega},y_{\omega})=\int_{0}^{t_{f}}\!\!\!\!\left<\begin{pmatrix}y_{\omega}(t)\\ u_{\omega}(t)\end{pmatrix},\begin{bmatrix}e^{\omega t}\mathcal{Q}(t)&e^{\omega t}\mathcal{N}^{*}(t)\\ e^{\omega t}\mathcal{N}(t)&e^{\omega t}\mathcal{R}(t)\end{bmatrix}\begin{pmatrix}y_{\omega}(t)\\ u_{\omega}(t)\end{pmatrix}\right>_{Y\times U}\!\!\!\!\!\!\!{\rm d}t$
@@
    along with boundedness of
    $C(sE-(A+BF))^{-1}B$
    on some half-plane. Namely, by invoking that, $(x,u,y)$ fulfills (1) if, and only if,
@@
    $J(u_{F},y_{F1},y_{F2})=\int_{0}^{t_{f}}\!\!\!\!\left<\left(\begin{smallmatrix}y_{F1}(t)\\ y_{F2}(t)\\ u_{F}(t)\end{smallmatrix}\right),\left[\begin{smallmatrix}\mathcal{Q}(t)&0&\mathcal{N}^{*}(t)\\ 0&0&-B\\ \mathcal{N}(t)&-B^{*}&\mathcal{R}(t)\end{smallmatrix}\right]\left(\begin{smallmatrix}y_{F1}(t)\\ y_{F2}(t)\\ u_{F}(t)\end{smallmatrix}\right)\right>_{Y^{2}\times U}\!\!\!\!\!\!\!{\rm d}t$
    subject to the above system. In the finite-dimensional case, the existence of a state feedback with such properties is guaranteed, if the system is impulse controllable, see [BR13]. The investigation of the existence of such feedbacks has not yet been conducted, to the best of the authors’ knowledge.

### v2

@@ 3 Exponential stability of abstract differential-algebraic equations
Utilizing the generator $A_{R}$ as described in (11), we can extend the characterization of exponential stability from Proposition 2.3 to pseudo-resolvents and degenerate semigroups. In this context, for a given degenerate semigroup $(T(t))_{t\geq 0}$ that is exponentially stable in the sense that it fulfills (13), we apply Laplace transform yielding a pseudo resolvent that is defined for all $\lambda\in\mathbb{C}$ with $\operatorname{Re}(\lambda)>\omega$, as discussed in [2], namely
@@
    There exists a positive operator $Q\in L(X)$ such that
    $\langle R(0)x,Qx\rangle+\langle Qx,R(0)x\rangle=-\langle R(0)x,R(0)x\rangle,\quad$
@@
    There exists a positive operator $Q\in L(X)$ such that
    $\langle R(0)x,Qx\rangle+\langle Qx,R(0)x\rangle\leq-\langle R(0)x,R(0)x\rangle,\quad$
@@
    $(-\omega,\infty)\subseteq\rho(E,A)$ and $(-\omega,\infty)\subseteq\rho(E^{*},A^{*})$;
@@
In particular, the degenerated semigroup associated with the pseudo-resolvent $\lambda\mapsto(A-\lambda E)^{-1}E$ is exponentially stable.
@@ 4 Linear quadratic optimal control of abstract differential-algebraic equations
In this section, we study LQ optimal control for ADAEs on a finite or infinite time horizons and we consider the pseudo-resolvent $R_{r}$ as in (8), and we assume that it fulfills the growth condition (9) for $\Omega=\{\lambda\in\mathbb{C}~{}|~{}\operatorname{Re}\lambda>\omega\}$ for some $\omega<0$. Further, let $T$ be the associated exponentially stable degenerated semi-group. With the notation as in Proposition 2.5, we use the bounded projection $P$ which fulfills ${\rm ran}\,P=X_{R}$ and $\ker P=X_{K}$ and decompose $A^{-1}Bu=(I-P)A^{-1}Bu+PA^{-1}Bu=x_{0}+x_{1}\in X$ and set $\hat{B}_{0}u:=(I-P)A^{-1}Bu$ and $\hat{B}_{1}u:=PA^{-1}Bu$ for all $u\in U$. Then $\hat{B}_{1}\in L(U,X_{R})$, $\hat{B}_{0}\in L(U,X_{K})$ and the mild solution of (5) is given by
@@
$\Psi:X\rightarrow L^{2}([0,t_{f}],Y),\quad(\Psi x_{0})(t):=CT(t)x_{0},\,\text{$t\in[0,t_{f}]$}.$
By invoking the boundedness assumption (4), we can, by using [19], infer that the input-output operator
$\mathbb{F}:\;$ $L^{2}([0,t_{f}],U)\rightarrow L^{2}([0,t_{f}],Y),$
@@
$y=\Psi x_{0}+\mathbb{F}u.$
In the following, we consider the time-varying weights $\mathcal{Q}(\cdot)\in L^{\infty}([0,t_{f}],\mathbb{R}^{n\times n})$, $\mathcal{R}(\cdot)\in L^{\infty}([0,t_{f}],\mathbb{R}^{m\times m})$, which are assumed to be symmetric almost everywhere and $\mathcal{N}(\cdot)\in L^{\infty}([0,t_{f}],\mathbb{R}^{m\times n})$ which are used in the quadratic cost function
$\hat{J}(u,y)$ $=\int_{0}^{t_{f}}\left\langle\begin{pmatrix}y(t)\\ u(t)\end{pmatrix},\begin{pmatrix}\mathcal{Q}(t)&\mathcal{N}^{*}(t)\\ \mathcal{N}(t)&\mathcal{R}(t)\end{pmatrix}\begin{pmatrix}y(t)\\ u(t)\end{pmatrix}\right\rangle_{Y\times U}{\rm d}t$
$=\left\langle\begin{pmatrix}y\\ u\end{pmatrix},\begin{pmatrix}\mathcal{Q}&\mathcal{N}^{*}\\ \mathcal{N}&\mathcal{R}\end{pmatrix}\begin{pmatrix}y\\ u\end{pmatrix}\right\rangle_{L^{2}}.$
@@
$J(u,x_{0}):=\hat{J}(u,\Psi x_{0}+\mathbb{F}u)$
$=\left\langle\begin{pmatrix}x_{0}\\ u\end{pmatrix},\begin{pmatrix}\Psi^{*}\mathcal{Q}\Psi&\Psi^{*}(\mathcal{Q}\mathbb{F}+\mathcal{N}^{*})\\ (\mathbb{F}^{*}\mathcal{Q}+\mathcal{N})\Psi&\mathscr{R}\end{pmatrix}\begin{pmatrix}x_{0}\\ u\end{pmatrix}\right\rangle$
@@
Assume that we have an abstract differential-algebraic system whose pseudo-resolvent $R_{r}$ is given by (8) and satisfies (9), (4) and assume that $\mathscr{R}$ satisfies the coercivity assumption (27). Then for all $x_{0}\in X$
@@
$u^{\rm opt}(t)=-\mathscr{R}^{-1}(\mathbb{F}^{*}\mathcal{Q}+\mathcal{N})\Psi(x_{0})(t),\,\text{for a.e.\ $t\geq 0$}.$ (28)
@@
We use an idea from [20], abbreviate $\hat{u}:=u+\mathscr{R}^{-1}(\mathbb{F}^{*}\mathcal{Q}+\mathcal{N})\Psi(x_{0})$ and complete the square as follows
@@
Since $\mathscr{R}$ is uniformly positive the minimum is attained if and only if the second term vanishes.
The formula (28) for the unique minimizer can be rewritten with $y(t)=\Psi(x_{0})(t)+\mathbb{F}u^{\rm opt}(t)$ as
@@
In addition to the assumptions of Proposition 4.1, let the cost function $J$ satisfy $\mathcal{N}=0$ and let there exists $\varepsilon>0$ such that $\langle\mathcal{R}u,u\rangle\geq\varepsilon\|u\|^{2}$ holds for all $u\in L^{2}([0,t_{f}],U)$. Then
$u^{\rm opt}(t)=-(I-\mathscr{R}^{-1}\mathbb{F}^{*}\mathcal{Q}\mathbb{F})^{-1}\mathscr{R}^{-1}\mathbb{F}^{*}\mathcal{Q}y(t).$ (29)
@@
$\|\mathscr{R}^{-1}\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|\leq\frac{\|\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|}{\|\mathcal{R}+\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|}\leq\frac{\|\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|}{\varepsilon+\|\mathbb{F}^{*}\mathcal{Q}\mathbb{F}\|}<1.$
Therefore the inverse of $I-\mathscr{R}^{-1}\mathbb{F}^{*}\mathcal{Q}\mathbb{F}$ can be obtained from the Neumann series.
If the coercivity condition (27) is violated then the existence of an optimal control input $u^{\rm opt}$ is not guaranteed, which can already be observed for finite-dimensional DAEs [17].
Remark 4.3 (Infinite Horizon).
@@
In particular, the result of Proposition 4.1 and Corollary 4.2 continues to hold. Well-posed linear systems that satisfy (30) with $E=I$ are called externally stable in the literature. For ODEs it was shown that for exponentially stabilizable and detectable systems the conditions (30) are equivalent to the exponential stability of the system (also called internal stability), see [13].
@@
The mild solutions that are considered here can be characterized using the mild solutions of the pseudo-resolvent equation (15) with $R=R_{r}$. It was shown in [7, Lemma 8.1] that one can also rewrite the ADAE (5) using the pseudo-resolvent $R=R_{l}=EA^{-1}$ in the following way
$\frac{{\rm d}}{{\rm d}t}EA^{-1}z(t)=z(t)+f(t)\quad EA^{-1}z(0)=EA^{-1}z_{0}.$ (32)
This rewriting does not change the inhomogeneity, but to guarantee the existence of solutions the initial value of the DAE is restricted to $x_{0}=A^{-1}z_{0}\in D(A)$.
@@
    $J(u_{\omega},y_{\omega})\\ =\int_{0}^{t_{f}}\!\!\!\!\left<\begin{pmatrix}y_{\omega}(t)\\ u_{\omega}(t)\end{pmatrix},\begin{bmatrix}e^{\omega t}\mathcal{Q}(t)&e^{\omega t}\mathcal{N}^{*}(t)\\ e^{\omega t}\mathcal{N}(t)&e^{\omega t}\mathcal{R}(t)\end{bmatrix}\begin{pmatrix}y_{\omega}(t)\\ u_{\omega}(t)\end{pmatrix}\right>_{Y\times U}\!\!\!\!\!\!\!{\rm d}t$
@@
    along with boundedness of $C(sE-(A+BF))^{-1}B$ on some half-plane. Namely, by invoking that, $(x,u,y)$ fulfills (1) if, and only if,
@@
    $J(u_{F},y_{F1},y_{F2})\\ =\int_{0}^{t_{f}}\!\!\!\!\left<\left(\begin{smallmatrix}y_{F1}(t)\\ y_{F2}(t)\\ u_{F}(t)\end{smallmatrix}\right),\left[\begin{smallmatrix}\mathcal{Q}(t)&0&\mathcal{N}^{*}(t)\\ 0&0&-B\\ \mathcal{N}(t)&-B^{*}&\mathcal{R}(t)\end{smallmatrix}\right]\left(\begin{smallmatrix}y_{F1}(t)\\ y_{F2}(t)\\ u_{F}(t)\end{smallmatrix}\right)\right>_{Y\times U^{2}}\!\!\!\!\!\!\!{\rm d}t$
    subject to the above system. In the finite-dimensional case, the existence of a state feedback with such properties is guaranteed, if the system is impulse controllable, see [4].
5 Application to heat-diffusion systems
A standard heat diffusion system on a one-dimensional spatial domain $[0,L]\subseteq\mathbb{R}$ with Dirichlet boundary conditions, is described by
$\frac{\partial T}{\partial t}(t,\xi)$ $=k\alpha\frac{\partial^{2}T}{\partial\xi^{2}}(t,\xi)+\mathbf{1}_{I_{U}}(\xi)u(t),$ (33)
$T(t,0)$ $=T(t,L)=0,\quad\text{for all $t\geq 0$,}$ (34)
where $T$ is the temperature, $\alpha>0$ is the diffusivity constant, $k>0$ is the thermal conductivity, $u\in L^{2}$ is a control input and $\mathbf{1}_{I_{U}}$ is the indicator function of a non-empty subinterval $I_{U}\subseteq[0,L]$. We choose $\mathcal{R}=I$, $\mathcal{Q}=I$, and $\mathcal{N}=0$, then $\mathscr{R}$ is coercive.
Following [11], we put (33) in the ADAE framework, by considering the defining relation between $T$ and the heat flux $J$ and Fourier’s law
$\frac{\partial T}{\partial t}=-\alpha\frac{\partial J}{\partial\zeta},\quad J=-k\frac{\partial T}{\partial\zeta}.$
Then the equivalent ADAE $\frac{\mathrm{d}}{\mathrm{d}t}Ex=Ax$ reads
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}\alpha^{-1}I&0\\ 0&0\end{bmatrix}\begin{bmatrix}T\\ k^{-1}J\end{bmatrix}$ $=\begin{bmatrix}0&-\tfrac{\partial}{\partial\zeta}\\ -\tfrac{\partial}{\partial\zeta}&-I\end{bmatrix}\begin{bmatrix}T\\ k^{-1}J\end{bmatrix}+\begin{bmatrix}0\\ \mathbf{1}_{I_{U}}\end{bmatrix}u(t)$
$y(t)$ $=\mathbf{1}_{I_{Y}}\alpha E\begin{bmatrix}T\\ k^{-1}J\end{bmatrix}=\mathbf{1}_{I_{Y}}(\cdot)T(t,\cdot)$
for some open interval $I_{Y}\subseteq[0,L]$ and with the operator $A$ that is given by
$A=\begin{bmatrix}0&\mathcal{D}^{*}\\ -\mathcal{D}&-I\end{bmatrix},\quad D(\mathcal{D})=\{x\in H^{1}([0,L])\,:\,x(0)=0\}.$
Hence, by [6, Lemma 2.7], the inverse $\lambda E-A$ exists for all $\lambda\in\mathbb{C}$ with $\operatorname{Re}\lambda\geq-\omega$ for some $\omega>0$ and the pseudo-resolvent $R_{r}$ is given by
$R_{r}(\lambda)$ $=(A-\lambda E)^{-1}E$
$=\begin{bmatrix}-I&\mathcal{D}^{*}\\ -\mathcal{D}&-\tfrac{\lambda}{\alpha}\end{bmatrix}\begin{bmatrix}(\tfrac{\lambda}{\alpha}+\mathcal{D}^{*}\mathcal{D})^{-1}&0\\ 0&(\tfrac{\lambda}{\alpha}+\mathcal{D}\mathcal{D}^{*})^{-1}\end{bmatrix}E$
$=\begin{bmatrix}-(\lambda+\alpha\mathcal{D}^{*}\mathcal{D})^{-1}&0\\ -\mathcal{D}(\lambda+\alpha\mathcal{D}^{*}\mathcal{D})^{-1}&0\end{bmatrix},$
where $\mathcal{D}^{*}=-\mathcal{D}$ and $\mathcal{D}\mathcal{D}^{*}=-\tfrac{\partial^{2}}{\partial\xi^{2}}$ with domain $H_{0}^{2}([0,L])$. Therefore, the resolvent growth assumption (9) and (4) hold. Therefore, the unique LQ optimal control is given by (29).